1.An assembly line consists of 3 tasks with individual times 5, 3, and 2 minutes and the cycle time of 5 minutes. Calculate the theoretical minimum number of workstations.

Select one:A. 3 B. 2 C. 6 D. 5 E. Cannot be calculated with the given information

2.Which is true for product layout?

Select one:A. Very flexible B. High job satisfaction C. Lower work-in-process inventories D. Easy to add new products

3.Which type of layout is best for high volumes of demand with few if any setups/changeovers?

Select one:A. Product B. Process C. Cellular D. Fixed position
Answer #1

1. Answer: B (2)

Rationale:

Total task time (TT) = 5+3+2 = 10 minutes

Cycle time (CT) = 5 minutes

Theoretical number of workstations = TT/CT

Theoretical number of workstations = 10/5

Theoretical number of workstations = 2

2. Answer: C (Lower work-in-process inventories)

Rationale: Product layouts generally have lower flexibility and therefore it is not easy to add new products. Product layouts have lower job satisfaction because of lower motivation as compared to process and fixed position layout. Some of advantages of product layout include lower cost, higher output and lower work-in-process inventories.

3. Answer: A (Product)

Rationale: Product layouts are commonly found in flow shops which produce high-volume and highly standardized products that require repetitive and highly standardized processes.